OpenStudy (anonymous):

Show that the equation x^5-3x^3=x+1 has a real root

OpenStudy (anonymous):

\[ x^5-3x^3=x+1 \\ x^5-3x^3-x-1=0\] polynomial is of odd degree goes from \(-\infty\) to \(\infty\) and is continuous, so takes on all values, including 0